OpenStudy (anonymous):

Convert the angle 13°38'35" to decimal degrees and round to the nearest hundredth

OpenStudy (freckles):

\[a^ob'c''=(a+\frac{b}{60}+\frac{c}{60^2})^o\]

OpenStudy (anonymous):

thank you…. is that the formula?

OpenStudy (anonymous):

13.63?

OpenStudy (freckles):

Yes because a minute is 60 parts of a degree and a second is 60 parts of a minute so a second is 3600 parts of a degree

OpenStudy (freckles):

I think I have 13.64 deg

OpenStudy (anonymous):

thank you

OpenStudy (freckles):

np